# Gravitational $D$-Form Factor: The $\sigma$-Meson as a Dilaton confronted with Lattice Data

### v1

@@ 3.2 Fit ansatz
In fact, it has been noted in the literature that $\sigma$-effects in the Euclidean domain are well approximated by a $\sigma$-pole with pole-mass in the range $500$-$600\,\mbox{MeV}$ [78] (see also [27, 29]), further supported by the empirical success of one-boson-exchange models (e.g. [79, 80]). The fact that the up and down quark masses are expected to enlarge the real part of the complex pole might well be compensated by the reduction of the imaginary because of reduced phase space. We therefore ignore this effect and adopt $m_{\mathrm{E},\sigma}=550(50)\,\mbox{MeV}$ as our central value. The parameter $r_{\mathrm{E},\sigma}$ serves as our primary fit variable, to be tested against the prediction of dilaton effective theory.
@@
The background contribution must of course be nonzero, as it accounts for higher states in the spectrum. In a spontaneously broken conformal theory, the dilaton (identified here with the $\sigma$-meson) is the only state that couples in the scalar channel, as follows from the textbook derivation of Goldstone’s theorem [81]. By contrast, the situation is less clear in a theory flowing to an infrared fixed point with spontaneous breaking of scale invariance. In such a scenario, suppression of other states near the fixed point appears plausible, since in the effective theory they would be loop-suppressed, unlike for the dilaton.
Since the above reasoning is not entirely rigorous it is best to take an agnostic attitude and explicitly parametrise for a background. For this purpose we use a combination of the following terms
$b(q^{2})=b+b^{\prime}q^{2}+\frac{r_{\text{eff}}}{q^{2}-m_{\text{eff}}^{2}}\;,$ (3.9)
@@
paying attention to avoiding over-parametrisation in view of the available data. Before presenting the fits, we digress to comment in more technical language why certain alternative parametrisations are not employed.
@@
    Are there viable alternatives to the Euclidean pole parametrisation? A first difficulty arises from the fact that, although the pole position itself is uniquely defined, the behaviour in its vicinity is not as it depends sensitively on the production mechanism or process under consideration. This issue is particularly relevant for the $\sigma$-meson, which is both broad (3.2) and located close to the left-hand cuts of $\pi\pi$-scattering. These features explain why the $\sigma$ phase shift does not pass rapidly through $180^{\circ}$; see, for example, Fig. 2 in [76] or the discussion in [82]. Through Watson’s theorem, this behaviour directly carries over to the form-factor case, since the same phase enters both $\pi\pi$-scattering and the form factor (via the Omnès representation [83]), see for instance [84, 85]. This property alone rules out many commonly used parametrisations, including all Breit–Wigner forms and Flatté-type (or “sill”) models [86]. While the sill model performs reasonably well for the moderately broad $a_{1}(1260)$ meson, it fails for the much broader $\kappa/K_{0}^{*}(700)$, which in fact shares important characteristics with the $\sigma$ [86]. Other approaches, such as the K-matrix employed by the HadSpec collaboration at $m_{\pi}\approx 239\,\mbox{MeV}$[87], or the single-resonance $S$-matrix solution of[82], are not suitable either, since they are designed for the $S$-matrix rather than for form factors. In scattering amplitudes, left-hand cuts are naturally included in these frameworks, which are absent in the form-factor case. A dispersive analysis using the Omnès representation [28, 30], is closer to extracting the complex-valued residue which is not what we are aiming at.
@@
    It has been argued that in the deep Euclidean limit $q^{2}\to$-$\infty$, the form factors scale as $D^{\pi}\propto 1/q^{2}$ and $D^{N}\propto 1/q^{6}$, based on light-cone sum rule computations [23, 24]; see also [88] for an instanton-based approach to the pion case. In the absence of further input, such asymptotic constraints can provide useful guidance for parametrisations [27, 29]. However, we choose not to impose them here for several reasons. First, it is not evident that $q^{2}=-2\,\mbox{GeV}^{2}$ should already be regarded as the deep Euclidean regime, and most of our relevant fit points are even less Euclidean. Second, the light-cone sum rule estimates rely heavily on the endpoint region of the corresponding distribution amplitudes, which are generally unknown and often assumed to take the asymptotic form. This assumption has been challenged by high-$q^{2}$ measurements of $\eta(\pi)\gamma\gamma^{*}$ transitions; see, for example, [89] and references therein.
@@ 4 Numerics and Fits to Lattice Data at $m_{\pi}\approx 170\,\mbox{MeV}$
In this section, we fit our ansatz (3.8) to lattice data at $m_{\pi}\approx 170\,\mbox{MeV}$ [16, 17], close to the physical pion mass of $m_{\pi}\approx 140\,\mbox{MeV}$. Our goal is to test whether the $\sigma$-meson ansatz naturally reproduces the $\sigma$-residue predicted by the dilaton interpretation Eq. (2.8). Attempting the converse, deriving the dilaton picture directly from the data, proves too challenging for the pion and only marginally feasible for the nucleon.
@@ 4.1 The nucleon gravitational form factor
$D^{N}(q^{2})=\frac{r_{\mathrm{E},\sigma}^{N}}{q^{2}-m_{\mathrm{E},\sigma}^{2}}+b+b^{{}^{\prime}}q^{2}+\frac{r_{\mathrm{eff}}}{q^{2}-m_{\mathrm{eff}}^{2}}\;,$ (4.1)
where we fix $m_{\mathrm{eff}}=1\,\mbox{GeV}$ to account for the effect of the $f_{0}(980)$ resonance and additional higher states in the spectrum. In our main fit we set $r_{\mathrm{eff}}=0$, while alternative combinations are used to test the robustness of the procedure. Since our primary goal is to compare the fit result with the dilaton prediction $r_{\sigma}^{N}=\frac{4}{3}\bar{m}_{N}^{2}$ (2.8), we need to estimate its central value and associated uncertainty. Here, $\bar{m}_{N}$ denotes the nucleon mass in the limit of vanishing $m_{u,d,s}$, for which we adopt the value $\bar{m}_{N}=826\,\mbox{MeV}$ from the recent pedagogical introduction [90]. [^5] Combining the uncertainties from the light and strange quark masses in quadrature gives a total uncertainty of roughly $10\,\mbox{MeV}$, which is small. One should, however, remain mindful of the longstanding tension between phenomenology and lattice results regarding light quark contributions [91]. This corresponds to a $13\%$ reduction from the physical nucleon mass, $m_{N}\approx 938\,\mbox{MeV}$, with a significant part arising from the nucleon’s strangeness content. As discussed earlier, the leading parametric corrections are radiative, of order ${\cal O}(m_{q}^{3/2})$, and are known to induce a negative shift of about $15\,\mbox{MeV}$ in the $SU(2)$ chiral perturbation theory [92]. Since the strange-quark contribution is comparable to that of the other light quarks, we assume a similar size for the radiative corrections and therefore double the uncertainty. In addition, one expects $q^{2}$-dependent corrections. For small $q^{2}$ these can be estimated within the effective theory, and their asymptotic behaviour suggests that they are not large (see the discussion in section 3.2). Such effects are absorbed into the background and are therefore, in principle, accounted for by the model uncertainty. To be conservative, we add an additional $10\%$ uncertainty, which combined in quadrature with the other effects yields the following estimate
@@
$\chi^{2}=\sum_{i,j=1}^{N_{\mathrm{data}}}(D^{N}_{\text{data}}-D^{N}_{\text{model}})_{i}\,(\text{Cov}^{-1})_{ij}\,(D^{N}_{\text{data}}-D^{N}_{\text{model}})_{j}\;,\quad\hat{\chi}^{2}\equiv\frac{\chi^{2}}{N_{\mathrm{dof}}}\;,$ (4.3)
using all 33 available data points [16] The effective number of degrees of freedom is $N_{\mathrm{dof}}=30$ since there are three fit parameters. A fit using the main parametrisation is displayed in the left figure of Fig. 1, where it is compared to the data of ref. [16]. Table 1 shows the corresponding fit parameters for a range of effective $\sigma$-masses, and the uncertainties correspond to the standard deviations as encoded in the covariance matrix. The covariance matrix of the fitted parameters is estimated from the inverse $\chi^{2}$-Hessian, evaluated at the $\chi^{2}$-minimum.
@@
Table 1: Model parameters for the nucleon $D$-form factor resulting from fits using our main background parametrisation, and for various values of the pole mass $m_{\mathrm{E},\sigma},$ to the $3$ lattice data points at $m_{\pi}\approx 170\,\mbox{MeV}$ [16]. The uncertainties correspond to the standard deviations encoded in the covariance matrix. The correlations between parameters are $(\rho_{rb},\,\rho_{rb^{\prime}},\,\rho_{bb^{\prime}})=(0.97,\,0.91,\,0.98)$ for $m_{\mathrm{E},\sigma}=550\,\mbox{MeV}$. LO dilaton effective theory predicts $r_{\sigma}^{N}=0.91(12)\,\mbox{GeV}^{2}$ (4.2).
@@
    $(r_{\mathrm{E},\sigma},b)^{N}$: with $r_{\mathrm{E},\sigma}^{N}=0.86(11)$, $b=0.291(77)$ and $\hat{\chi}^{2}=0.58$
@@
    $(r_{\mathrm{E},\sigma},b,r_{\mathrm{eff}})^{N}$: with $r_{\mathrm{E},\sigma}^{N}=1.55(58)$, b=0.04(28)b=-0.04(28), $r_{\mathrm{eff}}=-1.8(1.5)$ and $\hat{\chi}^{2}=0.55$
@@
    $(r_{\mathrm{E},\sigma},r_{\mathrm{eff}})^{N}$: with $r_{\mathrm{E},\sigma}^{N}=1.47(25)$, $r_{\mathrm{eff}}=-1.57(39)$ and $\hat{\chi}^{2}=0.53$
@@
    $(r_{\mathrm{E},\sigma},b^{\prime},r_{\mathrm{eff}})^{N}$: with $r_{\mathrm{E},\sigma}^{N}=1.53(45)$, $b^{\prime}=0.010(67)$, $r_{\mathrm{eff}}=-1.68(84)$ and $\hat{\chi}^{2}=0.55$
Firstly, we note that all of these fits are compatible with the main result (4.4). Secondly, they also remain consistent with the theoretical prediction (4.2), particularly once the spread in the effective mass is taken into account. Fits c) and d) provide a more direct test of the $\sigma$-meson dominance hypothesis, as the $r_{\mathrm{eff}}$-residue could potentially account for the curvature in the data. We observe that, while $r_{\mathrm{eff}}$ is non-negligible, $r_{\mathrm{E},\sigma}^{N}$ continues to agree with the theoretical expectation. In conclusion, there is good agreement within uncertainties with the dilaton interpretation, and this last observation suggests that the reverse scenario, establishing $\sigma$-dominance, is not out of sight.
Figure 1: The $D$-form factors are fitted using a Euclidean pole parametrisation (4.1) and (4.6) with $m_{\mathrm{E},\sigma}=550\,\mbox{MeV}$. The fits are compared to the lattice data, shown in black, for the nucleon [16] (left) and for the pion [17] (right). The dark curve indicates the central fit, while the shaded band represents the $68\%$ confidence interval.
Now, the complex-valued residue has been determined using Roy-Steiner equations [93]. Their result and our fit to the Euclidean residue differ, also in their absolute values

### v2

@@ 3.2 Fit ansatz
In fact, it has been noted in the literature that $\sigma$-effects in the Euclidean domain are well approximated by a $\sigma$-pole with pole-mass in the range $500$-$600\,\mbox{MeV}$ Donoghue:2006rg (see also RuizArriola:2010fj ; Broniowski:2024oyk ; Broniowski:2025ctl ), further supported by the empirical success of one-boson-exchange models (e.g. CalleCordon:2008eu ; Wu:2023uva ). The fact that the up and down quark masses are expected to enlarge the real part of the complex pole might well be compensated by the reduction of the imaginary part because of reduced phase space. We therefore ignore this effect and adopt $m_{\mathrm{E},\sigma}=550(50)\,\mbox{MeV}$ as our central value. The parameter $r_{\mathrm{E},\sigma}$ serves as our primary fit variable, to be tested against the prediction of dilaton effective theory.
@@
The background contribution must of course be nonzero, as it accounts for higher states in the spectrum. In a spontaneously broken conformal theory, the dilaton (identified here with the $\sigma$-meson) is the only state that couples in the scalar channel, as follows from the textbook derivation of Goldstone’s theorem Weinberg:1996kr . By contrast, the situation is less clear in a theory flowing to an infrared fixed point with spontaneous breaking of scale invariance. In such a scenario, suppression of other states near the fixed point appears plausible, since in the effective theory they would be loop-suppressed, unlike for the dilaton. We may gain insight by employing the dispersive techniques of Cao:2024zlf ; Cao:2025dkv , which are based on unitarity cuts and a two-channel Omnès solution, to infer the $D$-form factor (ultimately from experimental data and Roy–Steiner equations). For this purpose it is convenient to introduce a normalised trace of the energy-momentum tensor $\hat{\Theta}^{N}(q^{2})\equiv\Theta^{N}(q^{2})/(2m_{N}^{2})$, $\hat{\Theta}^{N}(0)=1$, and to consider the unsubtracted dispersion representation
$\Theta^{N}(q^{2})=\frac{1}{\pi}\int_{4m_{\pi}^{2}}^{\infty}\frac{ds\,\textrm{Im}\Theta^{N}(s)}{(s-q^{2}-i0)}\;.$ (3.9)
In order to make optimal use of data, the $\pi\pi$ and $K\bar{K}$ coupled-channel system is analysed in Cao:2024zlf ; Cao:2025dkv , together with an effective pole accounting for higher states, a strategy with a long-standing tradition, for instance in the description of electromagnetic form factors. To assess the $\sigma$-contribution, however, we restrict ourselves to the $\pi\pi$ channel, since the second resonance, $f_{0}(980)$, is known to predominantly couple to $K\bar{K}$. We therefore approximate the dispersive integral
$\Theta^{N}(q^{2})\approx\frac{1}{\pi}\int_{4m_{\pi}^{2}}^{4m_{K}^{2}}\frac{ds\,\textrm{Im}\Theta^{N}(s)}{(s-q^{2}-i0)}+\frac{c_{\Theta}}{1-q^{2}/m_{S}^{2}}\;,$ (3.10)
by cutting at the $K\bar{K}$ threshold, thus avoiding the coupled-channel region, and adding an effective pole that accounts for $f_{0}(980),f_{0}(1370)$ and further states. The single-channel imaginary part is given by Cao:2024zlf ; Cao:2025dkv
$\textrm{Im}\Theta^{N}(s)=-\frac{3\rho_{\pi}}{2s\rho_{N}^{2}}\left(f^{0}_{+}(s)\right)^{*}\Theta^{\pi}(s)\;,\quad\rho_{x}(s)\equiv\sqrt{1-4m_{x}^{2}/s}\;,$ (3.11)
where $\Theta^{\pi}(s)$ follows from the Omnès- and $f^{0}_{+}(s)$ from the $\pi N$ Roy–Steiner-solution. These quantities play the roles of $F_{\sigma}$ and $g_{\sigma NN}$ in (2.8).
The normalisation of the effective pole is fixed by the condition $\Theta^{N}(0)=1$, yielding $c_{\Theta}\approx-0.19(12)$ with uncertainties originating from the input to Eq. (3.11). [^5] We are indebted to the authors of Cao:2024zlf for sharing their data with us. The fact that the $\sigma$-meson contribution alone, within the one-channel approximation, saturates the sum rule reasonably well is remarkable and reminiscent of vector-meson dominance in the spin-$1$ channel, which itself still lacks a deeper explanation. It is therefore worth noting that, in the spontaneously broken scale invariance scenario, $\sigma$-dominance arises naturally, at least in the trace of the energy-momentum tensor.
Since we are fitting the $D$-form factor, its additional contributions must also be taken into account. These follow straightforwardly from the definition (2.1)
$D^{N}(q^{2})=\frac{4m_{N}^{2}}{3q^{2}}(A^{N}(q^{2})-\hat{\Theta}^{N}(q^{2}))-\frac{1}{3}(A^{N}(q^{2})-2J^{N}(q^{2}))\;.$ (3.12)
One observes that, in addition to the spin-0 channel, this expression involves the form factors $A$ and $J$, which are of spin-$2$ type, see for example Broniowski:2025ctl . The second term vanishes at $q^{2}=0$ and remains small throughout, whereas the $A$-form factor in the first term is sizeable, with $f_{2}(1270)$ being the first resonance in the spectrum.
We therefore conclude that, for the $D$-form factor, there is empirical evidence that the $\sigma$-meson is highly dominant in its spin-0 component, while the leading spin-$2$ resonance, $f_{2}(1270)$, appears at significantly higher mass. Moreover, we have verified that the $A$ (and $J$) form factors are very well described by a quadratic polynomial in $q^{2}$, which altogether motivates the parametrisation
$b(q^{2})=b+b^{\prime}q^{2}+b^{\prime\prime}q^{4}+\frac{r_{\text{eff}}}{q^{2}-m_{\text{eff}}^{2}}\;,$ (3.13)
with $m_{\text{eff}}=1.2\,\mbox{GeV}$ since our analysis indicates that $f_{2}(1270)$ is more prominent than $f_{0}(980)$. We note that (3.13) constitutes an over-parametrisation and would therefore lead to inflated uncertainties. Consequently, we restrict our fits to reduced parameter sets, typically involving three parameters including the $\sigma$ residue, and use the various fit combinations to assess robustness.
Finally, we emphasise that this background does not reflect a systematic effective-theory expansion in $q^{2}/(4\pi F_{\pi,\sigma})^{2}$ which would have a radius of convergence well below $2\,\mbox{GeV}^{2}$. Instead, the background, which also includes a constant term, effectively parametrises a sum of the most relevant higher resonances and multi-hadron states.
@@
Before presenting the fits, we digress to comment in more technical language why certain alternative parametrisations are not employed.
@@
    Are there viable alternatives to the Euclidean pole parametrisation? A first difficulty arises from the fact that, although the pole position itself is uniquely defined, the behaviour in its vicinity is not as it depends sensitively on the production mechanism or process under consideration. This issue is particularly relevant for the $\sigma$-meson, which is both broad (3.2) and located close to the left-hand cuts of $\pi\pi$-scattering. These features explain why the $\sigma$ phase shift does not pass rapidly through $180^{\circ}$; see, for example, Fig. 2 in Pelaez:2015qba or the discussion in Zheng:2003cr . Through Watson’s theorem, this behaviour directly carries over to the form-factor case, since the same phase enters both $\pi\pi$-scattering and the form factor (via the Omnès representation Omnes:1958hv ), see for instance Moussallam:2011zg ; Hoferichter:2012wf . This property alone rules out many commonly used parametrisations, including all Breit–Wigner forms and Flatté-type (or “sill”) models Giacosa:2021mbz . While the sill model performs reasonably well for the moderately broad $a_{1}(1260)$-meson, it fails for the much broader $\kappa/K_{0}^{*}(700)$, which in fact shares important characteristics with the $\sigma$ Giacosa:2021mbz . Other approaches, such as the K-matrix employed by the HadSpec collaboration at $m_{\pi}\approx 239\,\mbox{MeV}$ Rodas:2023nec , or the single-resonance $S$-matrix solution of Zheng:2003cr , are not suitable either, since they are designed for the $S$-matrix rather than for form factors. In scattering amplitudes, left-hand cuts are naturally included in these frameworks, which are absent in the form-factor case. A dispersive analysis using the Omnès representation Cao:2024zlf ; Cao:2025dkv , is closer to extracting the complex-valued residue which is not what we are aiming at.
@@
    It has been argued that in the deep Euclidean limit $q^{2}\to$-$\infty$, the form factors scale as $D^{\pi}\propto 1/q^{2}$ and $D^{N}\propto 1/q^{6}$, based on light-cone sum rule computations Tong:2021ctu ; Tong:2022zax ; see also Liu:2024vkj for an instanton-based approach to the pion case. In the absence of further input, such asymptotic constraints can provide useful guidance for parametrisations Broniowski:2024oyk ; Broniowski:2025ctl . However, we choose not to impose them here for several reasons. First, it is not evident that $q^{2}=-2\,\mbox{GeV}^{2}$ should already be regarded as the deep Euclidean regime, and most of our relevant fit points are even less Euclidean (see also the remarks in ref. Leutwyler:2002hm ). Second, the light-cone sum rule estimates rely heavily on the endpoint region of the corresponding distribution amplitudes, which are generally unknown and often assumed to take the asymptotic form. This assumption has been challenged by high-$q^{2}$ measurements of $\eta(\pi)\gamma\gamma^{*}$ transitions; see, for example, Agaev:2014wna and references therein.
@@ 4 Numerics and Fits to Lattice Data at $m_{\pi}\approx 170\,\mbox{MeV}$
In this section, we fit our ansatz (3.8) to lattice data at $m_{\pi}\approx 170\,\mbox{MeV}$ Hackett:2023rif ; Hackett:2023nkr , close to the physical pion mass of $m_{\pi}\approx 140\,\mbox{MeV}$. Our goal is to test whether the $\sigma$-meson ansatz naturally reproduces the $\sigma$-residue predicted by the dilaton interpretation Eq. (2.9). Attempting the converse, deriving the dilaton picture directly from the data, proves too challenging for the pion and only marginally feasible for the nucleon.
@@ 4.1 The nucleon gravitational form factor
$D^{N}(q^{2})=\frac{r_{\mathrm{E},\sigma}^{N}}{q^{2}-m_{\mathrm{E},\sigma}^{2}}+b+b^{{}^{\prime}}q^{2}+b^{{}^{\prime\prime}}q^{4}+\frac{r_{\mathrm{eff}}}{q^{2}-m_{\mathrm{eff}}^{2}}\;,$ (4.1)
where $m_{\mathrm{eff}}=1.2\,\mbox{GeV}$ as argued around Eq. (3.13). In our main fit we set $r_{\mathrm{eff}}=0$, while alternative combinations are used to test the robustness of the residue with respect to background variations. Since our primary goal is to compare the fit result with the dilaton prediction $r_{\sigma}^{N}=\frac{4}{3}\bar{m}_{N}^{2}$ (2.9), we need to estimate its central value and associated uncertainty. Here, $\bar{m}_{N}$ denotes the nucleon mass in the limit of vanishing $m_{u,d,s}$, for which we adopt the value $\bar{m}_{N}=826\,\mbox{MeV}$ from the recent pedagogical introduction Hoferichter:2025ubp . [^6] The uncertainty in $\bar{m}_{N}=826\,\mbox{MeV}$ is small but difficult to assess because of the longstanding tension between phenomenology and lattice results regarding light quark contributions, see FlavourLatticeAveragingGroupFLAG:2024oxs ; Hoferichter:2025ubp for references. A possible explanation could be excited state contamination Gupta:2021ahb . This corresponds to a $13\%$ reduction from the physical nucleon mass, $m_{N}\approx 938\,\mbox{MeV}$, with a significant part arising from the nucleon’s strangeness content. As discussed earlier, the leading parametric corrections to the nucleon residue, beyond known ${\cal O}(m_{q})$, are radiative corrections order of ${\cal O}(m_{q}^{3/2})$ (2.11). In order to estimate them we will take the nucleon mass corrections as a guidance. At $m_{\pi}\approx 170\,\mbox{MeV}$ the nucleon mass is expected to be around $970\,\mbox{MeV}$ as can be deduced from the mass decomposition Hoferichter:2025ubp or lattice plots for instance Owa:2023tbk . We will conservatively assign $50\%$ of this to the unknown ${\cal O}(m_{q}^{3/2})$ and higher corrections, which amounts to a $72\,\mbox{MeV}$-nucleon mass and $0.08\,\mbox{GeV}^{2}$-residue uncertainty, respectively. Let us return to the $q^{2}$-dependent vertex corrections mentioned at the end of section 3.1, which we estimate via the parametric dependence in the effective theory given by $q^{2}/(4\pi F_{\sigma})^{2}$. From $F_{\sigma}^{2}\propto N_{f}$ and since $F_{\pi}=F_{\sigma}$ in the $N_{f}=2$ linear $\sigma$-model, we take $F_{\sigma}^{2}=3/2,F_{\pi}^{2}$. The pion decay constant at $m_{\pi}\approx 170\,\mbox{MeV}$ is $F_{\pi}\approx 100\,\mbox{MeV}$, as inferred from the plots in Brandt:2013dua . Since $q^{2}/(q^{2}-m_{\mathrm{E},\sigma}^{2})=m_{\mathrm{E},\sigma}^{2}/(q^{2}-m_{\mathrm{E},\sigma}^{2})+1$ and the constant term can be absorbed into the background, this implies a relative uncertainty of $m_{\mathrm{E},\sigma}^{2}/(4\pi F_{\sigma})^{2}\approx 0.13$. Adding both uncertainties in quadrature yields the estimate
@@
$\chi^{2}=\sum_{i,j=1}^{N_{\mathrm{data}}}(D^{N}_{\text{data}}-D^{N}_{\text{model}})_{i}\,(\text{Cov}^{-1})_{ij}\,(D^{N}_{\text{data}}-D^{N}_{\text{model}})_{j}\;,\qquad\hat{\chi}^{2}\equiv\frac{\chi^{2}}{N_{\mathrm{dof}}}\;,$ (4.3)
using all 33 available data points Hackett:2023rif . The effective number of degrees of freedom is $N_{\mathrm{dof}}=30$ since there are three fit parameters. A fit using the main parametrisation is displayed in the left figure of Fig. 1, where it is compared to the data of ref. Hackett:2023rif . Table 1 shows the corresponding fit parameters for a range of effective $\sigma$-masses, and the uncertainties correspond to the standard deviations as encoded in the covariance matrix. The covariance matrix of the fitted parameters is estimated from the inverse $\chi^{2}$-Hessian, evaluated at the $\chi^{2}$-minimum.
@@
Table 1: Model parameters for the nucleon $D$-form factor resulting from fits using our main background parametrisation, and for various values of the pole mass $m_{\mathrm{E},\sigma},$ to the $3$ lattice data points at $m_{\pi}\approx 170\,\mbox{MeV}$ Hackett:2023rif . The uncertainties correspond to the standard deviations encoded in the covariance matrix. The correlations between parameters are $(\rho_{rb},\,\rho_{rb^{\prime}},\,\rho_{bb^{\prime}})=(0.97,\,0.91,\,0.98)$ for $m_{\mathrm{E},\sigma}=550\,\mbox{MeV}$. LO dilaton effective theory predicts $r_{\sigma}^{N}=0.91(14)\,\mbox{GeV}^{2}$ (4.2).
@@
    $\{r_{\mathrm{E},\sigma},b\}^{N}=\{\,0.86(11)\,,\,0.291(77)\}$ and $\hat{\chi}^{2}=0.58$
@@
    $\{r_{\mathrm{E},\sigma},b,r_{\mathrm{eff}}\}^{N}=\{\,1.40(46)\,,\,-0.13(35)\,,\,-2.1(1.7)\}$ and $\hat{\chi}^{2}=0.55$
@@
    $\{r_{\mathrm{E},\sigma},r_{\mathrm{eff}}\}^{N}=\{\,1.25(20)\,,\,-1.49(38)\}$ and $\hat{\chi}^{2}=0.54$
@@
    $\{r_{\mathrm{E},\sigma},b^{\prime},r_{\mathrm{eff}}\}^{N}=\{\,1.36(36)\,,\,0.026(74)\,,\,-1.78(89)\}$ and $\hat{\chi}^{2}=0.55$
  • e)
    $\{r_{\mathrm{E},\sigma},b,b^{\prime},b^{\prime\prime}\}^{N}=\{\,1.37(57)\,,\,1.2(1.1)\,,\,0.026(74)\,,\,0.12(26)\}$ and $\hat{\chi}^{2}=0.57$
Firstly, we note that all of these fits are compatible with the main result (4.4). Secondly, they also remain consistent with the theoretical prediction (4.2), particularly once the spread in the effective mass is taken into account. Fits c) and d) provide a more direct test of the $\sigma$-meson dominance hypothesis, as the $r_{\mathrm{eff}}$-residue could potentially account for the curvature in the data. We observe that, while $r_{\mathrm{eff}}$ is non-negligible, $r_{\mathrm{E},\sigma}^{N}$ continues to agree with the theoretical expectation. We further note that reducing the fit interval to $0,-1.5]\,\mbox{GeV}^{2}$ and $0,-1]\,\mbox{GeV}^{2}$ for the linear background leads to $\{r_{\mathrm{E},\sigma}^{N},\hat{\chi}^{2}\}=\{1.19(34),0.6\}$ and $\{r_{\mathrm{E},\sigma}^{N},\hat{\chi}^{2}\}=\{1.34(72),0.46\}$, respectively, which further hints at the robustness of the result. In conclusion, there is good agreement within uncertainties with the dilaton interpretation, and this last observation suggests that the reverse scenario, establishing $\sigma$-dominance, is not out of sight.
Figure 1: The $D$-form factors are fitted using a Euclidean pole parametrisation (4.1) and (4.6) with $r_{\mathrm{eff}}=0$ and $m_{\mathrm{E},\sigma}=550\,\mbox{MeV}$. The fits are compared to the lattice data, shown in black, for the nucleon Hackett:2023rif (left) and for the pion Hackett:2023nkr (right). The dark curve indicates the central fit, while the shaded band represents the $68\%$ confidence interval.
Now, the complex-valued residue has been determined using Roy-Steiner equations Hoferichter:2023mgy . Their result and our fit to the Euclidean residue differ, also in their absolute values